# A Comprehensive Study on GDPR-Oriented Analysis of Privacy Policies: Taxonomy, Corpus and GDPR Concept Classifiers

### v1

@@ IV-B4 Constructing Hierarchical Structure
Step 1: Pre-processing. Although many HTML elements have been removed as part of the first pre-processing step described in Section IV-B2, some ad hoc inline elements, such as those appearing in the middle of a <p> element, play no roles in PP-XML. We remove such inline elements. At the end, we put all remaining elements into a <policy> element.
@@
Step 2: Initial conversion of some HTML elements to PP-XML elements. As shown in Table III, there are a precise mapping from some HTML elements to <list> and <item> elements in PP-XML, so such HTML elements can be directly converted.
@@
TABLE III: Different types of PP-XML elements and HTML elements that may be converted to each PP-XML element type
@@
<policy> The privacy policy as a whole (the root element) NA (determined based on the method introduced in Section IV-B3)
<segment> A semantic segment of the privacy policy (child of the <policy> element or another <segment> element) NA (derived from <title> elements, see Step 4 discussed in Section IV-B4)
@@
Step 3: Classifying the remaining HTML elements into title elements at different levels and paragraph elements in PP-XML. In privacy policies, titles tend to have different attributes from paragraphs, e.g., titles are often in bold face and have a shorter text length. Based on manual inspection of title and paragraph elements of some randomly selected privacy policies, we identified the following attributes that may be useful for differentiating title elements from paragraph elements: i) the text length, ii) the font size, iii) the font weight, iv) if the text is italic, v) if the text is underlined, vi) the node depth in the HTML DOM tree, vii) the HTML tag, and viii) leading ordinal labels (LOLs) for titles (e.g., ‘1’, ‘2.3’, ‘a.’, and ‘ii)’) (see Table IV for more examples on LOLs). As part of the input features of the title and paragraph classifier, a 12-D descriptor is used to describe the leading ordinal label of a candidate title or paragraph element. The 12-D vector is composed of four 3-D sub-vectors, each representing a sub-label of the leading ordinal label at one of the four possible title levels. Each sub-vector follows the format (sub-label’s format, sub-label’s value, sub-separator’s format). The sub-label’s format and the sub-separator’s format are determined according to Table IV. The sub-label’s value is the 1-based natural ordinal value of the label, e.g., the number itself for Arabic numbers, 1 for ‘a’, ‘A’ and ‘i’. For instance, the sub-label ‘3.’ will be mapped to (1, 3, 1), ‘b)’ to (3, 2, 3), and a full label ‘3.a.i’ will be mapped to the 12-D descriptor [1 3 1 2 1 1 4 1 0 0 0 0].
In total, for each candidate title/paragraph element, we have these input features and an ML-based classifier can be constructed to predict the candidate element into one of the following five classes: title at Level $i$ ($i=1,2,3,4$), and paragraphs. We tested several mainstream ML models that do not require a large training set for the multi-class classification task, including random forest (RF) [56], SVM (with linear kernel and RBF kernel) [57], Extra Trees (ET) [58], and XGBoost [59]. We used all the 150 privacy policies for training and testing purposes. We used 5-fold cross validation and 20% as the testing set. Table V shows the results of all classifiers, indicating that the ET classifier achieved the best performance, with the F1-score reaching 0.882. As can be seen, the performance is generally good enough to support the semi-automated corpus construction process.
TABLE IV: Mappings between sub-label’s format and sub-separator’s format and values in the leading ordinal label descriptor
@@ 5 Others Others
TABLE V: Results of title and paragraph classifiers
@@
IV-B5 GDPR Concept Annotation
Using the process described in Section IV-B4, we processed all 150 privacy policies in our corpus to get 150 PP-XML documents. Then, we annotated the 150 PP-XML documents with relevant GDPR concepts based on our extended GDPR taxonomy introduced in Section IV-A. The annotation was done for each title and paragraph element in each PP-XML document, with one or more tags each representing a unique GDPR node in our extended GDPR taxonomy.
For the annotation work, we decided to pay a commercial annotation company to ensure the quality of the work. We followed a multi-stepped quality assurance process. In Step 1, one of the authors of this paper, who has good knowledge of GDPR concepts, first trained four annotators of the company, and then the annotators attempted to annotate 20 PP-XML documents as a pilot. In Step 2, the results of the pilot annotation experiments were checked by the author and feedback was given to the four human annotators so they knew how to refine and align their annotations. In Step 3, the four annotators went ahead to annotate all the remaining PP-XML documents so that each document was annotated independently by two different annotators. In Step 4, a random sample of the annotation results of each annotator was checked by another annotator. Finally, the full annotation results were checked by three authors of this paper, who then discussed the results between them to agree on any different annotations. To quantitatively measure the quality of the four annotators’ work, we used the inter-rater agreement of Cohen’s Kappa ($\kappa$[60], and observed an average $\kappa$ score of 0.75 across all the annotated documents, indicating that the two independent human annotators had a substantial level of agreement for most documents. For disagreements, we observed that most are about the GDPR concepts ‘PD CATEGORY’, ‘PD ORIGIN’, and ‘PROCESSING PURPOSES’. The reason is that these concepts tend to be declared together in many privacy policies, making it difficult to separate them. At the end of the annotation work, we finally obtained a fully annotated privacy policy corpus GoPPC-150 following PP-XML and our extended GDPR taxonomy.
V GDPR Concept Classifiers and Evaluation Methods
As mentioned in Section II, past studies on automated GDPR compliance of privacy policies focuses mostly on long texts or paragraphs in the privacy policy without considering the hierarchical structure, e.g., no previous work has considered automatic analysis of titles in a privacy policy. Missing such context information in a privacy policy can lead to an incomplete and less accurate analysis. What’s more, previous work tends to view compliance checking task as multi-label classification, using one model classify all concepts of the same level. As a result, the input features of the model are all the same. This seems work when concept types are few. But for the 96 concepts of different levels in our GDPR taxonomy, the same features may well are not suitable. To figure out the effect of different input features and structures of models, we conducted a comprehensive study. As shown in Figure 3, we tried different combinations of structures of classifiers, input features, corpora and evaluation methods, then compared each performance on each concept.
Noticeably, the binary RF model with 100-D TF-IDF vectorization input features is enough for titles’ classification, and the length of titles’ text limit the experiment much. So we only conducted the comprehensive study on paragraphs and applied the simple configuration on title classification task.
Figure 3: Comprehensive study of concept classifiers
V-A Structures of Classifiers
As mentioned in Section II, compliance checking of privacy policy can be regard as an HMTC task and performed in three ways: flat, local and global approaches. Within local approaches [8], there are LCN, LCPN, LCL and so on [10, 11, 12]. Until now, related work has focused on the LCPN approach, using one classifier classifying all concepts of the same level. But we also tried LCN approach as it is flexible to set different configurations for different concepts. When there are many concepts to be classified, LCN classifiers allow us to train the best models for specific concepts. When the samples of specific concepts are not enough, upsampling can be applied without affecting other concepts’ samples in LCN classifiers. So we tried LCPN and LCN approaches to construct the classifiers. All systematic experiments are both performed on two classifier types respectively.
V-B Input Features
When classifying a paragraph or title in a privacy policy, we considered the following three sub-groups of features as the input of the classifier: (1) the vector embeddings of the paragraph or title, (2) the vector embeddings of the already processed parent and sibling node of the paragraph or title, and (3) a binary vector representing if some pre-defined keywords associated with each GDPR concept appears or not in the current paragraph or title. The first-subgroup of features cover all vector embeddings used in past studies, and the third one was proposed by Amaral et al. [25]. The second subgroup is new features proposed by us for this work to cover capture the hierarchical information between the current node and its parent and the contextual information between the current node and its siblings under the same parent.
For the first sub-group of features on vector embeddings, we first tested three widely used traditional vectorization methods with three different dimensionalities (100-D, 200-D and 300-D): TF-IDF [61], GloVe [62], fastText [63]. We also tested combinations of three different pairs of the three traditional vectorization methods, with a 100-D subvector for each method and a 200-D vector for the combination. The experimental results showed that 300-D TF-IDF performed the best for binary classifiers and we implemented the methods as the default vectorization for binary classifiers. We then compared TF-IDF and the more advanced vectorization method based on PrivBERT [23] as two alternative choices for our more comprehensive experiments.
For the second sub-group of features on capturing hierarchical and contextual information, we use the vector embeddings of already processed parent and sibling nodes of the current node in the same privacy policy. Considering parent nodes tend to be titles while sibling nodes tend to be paragraphs when the current node is a paragraph, we considered that 100-D is enough for the former and 300-D is suitable for the latter. The vector embedding methods considered were also TF-IDF and PrivBERT. These features help us capture the hierarchical information between the current node and its parent and sibling nodes. Our experiments prove the features improved the performance of classifiers as shown in Tables VII and VIII.
For the third sub-groups of features about occurrences of pre-defined keywords, we used a 96-D binary vector (the number of all nodes in our GDPR taxonomy), where each feature indicates if at least one pre-defined keywords associated with a specific GDPR concept appears in the candidate element at least once. For example, a sentence like “When Google shares your information” will be labeled as ‘DATA SHARING.CONDITION’ because it contains the keywords ‘share’ and ‘when’. Because Amaral et al. [25] did not release their pre-defined keywords and our GDPR taxonomy includes more concepts, we decided to define our own keywords list for each of the 96 concepts. The full list of such keywords can be found in the supplementary data file available at https://github.com/tp-sh/GDPR_privacy_policies/blob/main/data/keyword_list.csv). Our experimental results in Tables VII and VIII show that these keyword-based features could help improve the performance of LCN classifiers.
For the LCPN classifiers, we set two sub-groups of input features, which are PrivBERT embeddings [23] of current node’s text and PrivBERT embeddings of the parent’s and sibling’s text’s embeddings. The experiments showed the parent’s and sibling’s embeddings had few help in classifications, so we used the PrivBERT embeddings of current node’s text as the input features of the fine-tuned PrivBERT.
V-C Corpus
As the OPP-115 corpus is labeled at the segment-level, it cannot accomplish document-level evaluation. Therefore, we use our corpus GoPPC-150 as the corpus of classifiers, which contains manual annotations of 14K (8k paragraphs and 6k titles) fine-grained data practices on 150 privacy policies. The annotations contain 96 nodes of three levels.
V-D ML Models
For LCN classification, Torre et al. [7] reported that the SVM model achieved the best performance. Therefore, we adopted the SVM model for our testing and conducted comparison experiments using other mainstream ML models, including random forest (RF), XGBoost, ET and neural network (MLP). We used 300-D TF-IDF features of current node as input to test the performance of these models, and the results showed that RF achieved the best performance for the most concept. The full results of all classifiers can be found at https://github.com/tp-sh/GDPR_privacy_policies/blob/main/data/comparison_experiments.csv. Therefore, we took RF classifiers as the representative of LCN classifiers. For LCPN classification, a multi-class classifier is required to classify all the child nodes under the same parent node. In the current literature, most studies [20, 21, 22, 23] use multi-class neural network models as the classifier. Therefore, we opted for multi-class neural network as the representative of LCPN classifiers.
V-E Evaluation Frameworks
Different from evaluating classifier performance at the segment level in the literature, we propose a document-level evaluation framework that is more suitable for real-world scenarios. A detailed comparison of segment-level and document-level evaluation frameworks is shown in Figure 4. Document-level evaluation is achieved by selecting some documents in the corpus to form the training set to train the classifier, and using the remaining documents as the test set. In the testing phase, the trained classifiers are used to make predictions for different concepts for each paragraph and title in each document of the test set, and the performance metrics of each classifier are determined according to the predicted results. The advantage of this evaluation framework is that it is more in line with real-world application scenarios, that is, the trained classifiers are applied to process new (i.e., unseen) privacy policies. For our task, we chose 120 privacy policies in our corpus GoPPC-150 as the training set, and the rest 30 privacy policies as the test set.
Figure 4: Evaluation methods
By trying different combinations of input features, classifier structures and models, we got 12 types of classifiers (see Table VI for the detailed configurations of the 12 classifiers) and the results of each classifier. We used three common metrics, precision (P), recall (R), and F1-scores, to evaluate the performance of each classifier. The difference of performance between each two types can show the effect of the different configurations. The overall performances of selected classifiers are shown in Table VII and detailed performances of some examples are shown in Table VIII. Due to the space limit, we only show the F1-scores of representative 6 classifiers in the two tables. The full results of all classifiers can be found at https://github.com/tp-sh/GDPR_privacy_policies/blob/main/data/complete_classifier_results.csv.
TABLE VI: The 12 different types of classifiers
Type Input Features Machine Learning Model
1 300-D TF-IDF features of the current node Binary RF
2 300-D TF-IDF features of the current, parent and sibling nodes Binary RF
3 300-D TF-IDF features of the current node, keyword occurrences Binary RF
4 300-D TF-IDF features of the current, parent and sibling nodes, keyword occurrences Binary RF
5 PrivBERT embeddings of the current node Binary RF
6 PrivBERT embeddings of the current, parent and sibling nodes Binary RF
7 300-D TF-IDF features of the current node Multi-class neural network
8 300-D TF-IDF features of the current, parent and sibling nodes Multi-class neural network
9 300-D TF-IDF features of the current node, keyword occurrences Multi-class neural network
10 300-D TF-IDF features of the current, parent and sibling nodes, keyword occurrences Multi-class neural network
11 PrivBERT embeddings of the current node Multi-class neural network
12 PrivBERT embeddings of the current, parent and sibling nodes Multi-class neural network
@@
TABLE VII: Overall performance results in both segment-level and document-level evaluation frameworks (F1-scores)
Evaluation Segment-level Document-level
Macro averages of level-1 nodes 0.748 0.797 0.753 0.804 0.773 0.805 0.512 0.558 0.509 0.554 0.658 0.669
Macro averages of all nodes 0.702 0.717 0.702 0.716 0.666 0.682 0.453 0.454 0.426 0.448 0.527 0.529
Ratio of maximum F1-score in level-1 nodes 0/14 0/14 1/14 0/14 0/14 1/14 0/14 1/14 0/14 1/14 4/14 3/14
Ratio of maximum F1-score in all nodes 4/35 4/35 3/35 1/35 0/35 1/35 3/35 2/35 1/35 2/35 8/35 5/35
TABLE VIII: Detail performance results (F1-scores). For each cell, the performance figure for the segment-level evaluation is shown first, followed by that of the document-level evaluation in parentheses. For LCPN classifiers (Types 11 and 12) on Level 2, their performances are based on the those of two cascaded classifiers on Levels 1 and 2. In the Name column, DSR represents “DATA SUBJECT RIGHT”, DS represents “DATA SHARING”, LB represents “LAWFUL BASIS”.
Concept Type 1 Type 2 Type 3 Type 4 Type 11 Type 12
Level Name
1 CONTROLLER 0.749 (0.519) 0.726 (0.554) 0.788 (0.598) 0.738 (0.455) 0.741 (0.593) 0.735 (0.575)
DATA SUBJECT RIGHT 0.783 (0.493) 0.833 (0.477) 0.799 (0.350) 0.821 (0.497) 0.816 (0.658) 0.828 (0.694)
PD ORIGIN 0.775 (0.538) 0.852 (0.655) 0.779 (0.444) 0.845 (0.699) 0.825 (0.668) 0.842 (0.770)
PD SECURITY 0.752 (0.792) 0.857 (0.698) 0.777 (0.737) 0.844 (0.846) 0.800 (0.767) 0.840 (0.744)
DATA SHARING 0.764 (0.640) 0.833 (0.697) 0.761 (0.716) 0.837 (0.764) 0.806 (0.757) 0.849 (0.772)
2 DSR.RESTRICTION 0.778 (0.457) 0.694 (0.400) 0.781 (0.359) 0.658 (0.451) 0.733 (0.557) 0.731 (0.556)
DSR.OBJECT 0.684 (0.440) 0.747 (0.190) 0.675 (0.237) 0.628 (0.172) 0.586 (0.382) 0.512 (0.425)
LB.LEGITIMATE INTEREST 0.729 (0.419) 0.564 (0.280) 0.667 (0.421) 0.709 (0.235) 0.536 (0.426) 0.540 (0.440)
LB.CONSENT 0.424 (0.244) 0.313 (0.098) 0.486 (0.195) 0.389 (0.108) 0.333 (0.244) 0.308 (0.280)
DS.RECIPIENT 0.637 (0.562) 0.713 (0.570) 0.618 (0.537) 0.715 (0.626) 0.639 (0.516) 0.681 (0.554)
Table VII shows that, for some GDPR concepts, non-PrivBERT classifiers could outperform PrivBERT-based ones with the segment-level evaluation method. Our results give a very different picture from what was reported in [23], which showed PrivBERT-based classifiers as the best. Despite differences under the document-level evaluation framework, PrivBERT-based classifiers performed better in comparison with those non-PrivBERT ones, but all classifiers showed significant declines in F1 score, which were not sufficient to support automation analysis of privacy policies. An important insight learned here is that more comprehensive comparative studies across multiple settings like what we did are needed to evaluate performances of GDPR concept classifiers. What’s more, under the document-level evaluation framework, PrivBERT-based classifiers have the best macro average F1-scores for all nodes, it shows the best performance in most nodes. However, the results in the segment-level evaluation case are less stable, as the best performance appears under different types for different concepts. This indicates that the document-level evaluation framework may be able to produce more consistent results for deciding the best feature set.
The detail results in Table VIII show that the effects of different types of features are different for different concept classifiers, for both segment-level and document-level evaluation frameworks. The results were expected by us because unique characteristics of different concepts may be better captured by different combinations of features. For instance, for the contextual features we introduced in this paper, this concept-dependency may be explained as follows: some concepts are ambiguous and contextual features can help distinguish them from others, but for some other concepts contextual features may just add more redundant information and confuse the learning process. As a result, the optimal input features of different GDPR concept classifiers may need to be different. This strategy can be more easily applied to LCN classifiers than to LCPN classifiers since the latter combine multiple concepts. The structure of LCPN classifiers also means inevitable cascade errors. As shown in Table VIII, when it comes to level-2 nodes, the LCPN classifiers’ cascade performances go down due to the limit of classifiers at level 1. What’s more, all concepts at the same level need to be trained in LCPN classifications even if there are not enough data for one or more concepts involved, while for LCN classifiers we can choose not to train a specific classifier if there are not enough training samples.
VI Further discussion
The significant difference between the performance results obtained by these two evaluation methods further demonstrates the value of document-level evaluation methods. The results obtained by the previous segment-level evaluation methods are somewhat inflated, and the excellent performance in the test does not mean that the correct prediction can be made in the real environment. Only by using the document-level evaluation method can we understand the performance of the classifiers in the real environment. However, previous work did not pay attention to this point, which may be because the current open source corpus is also segment-level, such as our document-divided corpus is not common, so relevant work can only use segment-level assessment methods. Therefore, we propose that the future corpus need to be divided into documents, so that the document-level evaluation method can be adopted.
The results of our evaluation method show that the performance of current concept classifiers is overestimated, and in fact, the performance of classifiers can not meet the requirements of large-scale automated detection. This results in an error bar that is too large to be trusted. Therefore, it is not possible to do large-scale automated detection of privacy policies at present, and the focus of current work should also be to improve the performance of classifiers. In order to achieve this goal, the corpus needs to be greatly expanded, which cannot be done by our efforts alone, and requires the cooperation of the entire community. We plan to set up an open website and open source our corpus and code to facilitate such a community-wide effort.
At the same time, we found that a single set of classifiers can not always obtain the best performance in all concepts, and for some concepts, the simple configuration of classifiers can get better performance, which indicates that the previous ”one size fits all” point is not correct, especially for the privacy policy analysis scenario that requires dozens of classifiers. It is obviously far-fetched to always adopt the same configuration. For obtaining the best classifier for each concept, it is necessary to conduct comprehensive experiments and analysis to get the best classifier configuration.
Furthermore, LLM may also represent a potential development direction of this part of work. The large training samples of LLM can mitigate the limitations of corpora and ensure the robustness of the classification.
VII Conclusion
This paper proposes a new evaluation method for concept classifiers which can present the performance of classifiers in real environment. Different from previous method, the new approach divides samples in document-level while training and evaluating. The gap between our evaluation results and previous work shows that the former classifiers are highly overestimated. The real performance of the classifiers are not enough for the large scale automation compliance analysis. We also noticed some other gaps in previous work and improved them. This includes the most complete GDPR-oriented privacy policy taxonomy, a new GDPR-oriented privacy policy corpus GoPPC-150 and an automation-enhancing frameworks for supporting future research on GDPR-oriented analysis of privacy policies. This paper also provides the most comprehensive study on the GDPR-oriented analysis of privacy policies. The study led to a range of new findings and insights, including the usefulness of context-based features for improving performance of machine learning based GDPR concept classifiers, the importance to consider the LCN architecture of hierarchical classifiers and testing cross-corpus generalizability, and the finding that a “one size fits all” idea does not work for GDPR concept classifiers so a more locally optimized approach is needed. Our study also led to a range of new GDPR concept classifiers that can be used by researchers and practitioners. Although our work focused on GDPR-oriented analysis, most results and outcomes can either be easily extended to study other and multiple data protection laws, or the general insights can be re-validated in the wider data protection landscape.

### v2

@@ Appendix B The process of constructing hierarchical structure
Step 1: Pre-processing. Although many HTML elements have been removed as part of the first pre-processing step described in Section III-B2, some ad hoc inline elements, such as those appearing in the middle of a <p> element, play no roles in PP-XML. We remove such inline elements. At the end, we put all remaining elements into a <policy> element.
@@
Step 2: Initial conversion of some HTML elements to PP-XML elements. As shown in Table V, there are a precise mapping from some HTML elements to <list> and <item> elements in PP-XML, so such HTML elements can be directly converted.
@@
TABLE V: Different types of $P$-XML elements and HTML elements that may be converted to each PP-XML element type
@@
<policy> The privacy policy as a whole (the root element) NA (determined based on the method introduced in Section III-B3)
<segment> A semantic segment of the privacy policy (child of the <policy> element or another <segment> element) NA (derived from <title> elements, see Step 4 discussed in Section III-B4)
@@
Step 3: Classifying the remaining HTML elements into title elements at different levels and paragraph elements in PP-XML. In privacy policies, titles tend to have different attributes from paragraphs, e.g., titles are often in bold face and have a shorter text length. Based on manual inspection of title and paragraph elements of some randomly selected privacy policies, we identified the following attributes that may be useful for differentiating title elements from paragraph elements: i) the text length, ii) the font size, iii) the font weight, iv) if the text is italic, v) if the text is underlined, vi) the node depth in the HTML DOM tree, vii) the HTML tag, and viii) leading ordinal labels (LOLs) for titles (e.g., ‘1’, ‘2.3’, ‘a.’, and ‘ii)’) (see Table VI for more examples on LOLs). As part of the input features of the title and paragraph classifier, a 12-D descriptor is used to describe the leading ordinal label of a candidate title or paragraph element. The 12-D vector is composed of four 3-D sub-vectors, each representing a sub-label of the leading ordinal label at one of the four possible title levels. Each sub-vector follows the format (sub-label’s format, sub-label’s value, sub-separator’s format). The sub-label’s format and the sub-separator’s format are determined according to Table VI. The sub-label’s value is the 1-based natural ordinal value of the label, e.g., the number itself for Arabic numbers, 1 for ‘a’, ‘A’ and ‘i’. For instance, the sub-label ‘3.’ will be mapped to (1, 3, 1), ‘b)’ to (3, 2, 3), and a full label ‘3.a.i’ will be mapped to the 12-D descriptor [1 3 1 2 1 1 4 1 0 0 0 0].
In total, for each candidate title/paragraph element, we have these input features and an ML-based classifier can be constructed to predict the candidate element into one of the following five classes: title at Level $i$ ($i=1,2,3,4$), and paragraphs. We tested several mainstream ML models that do not require a large training set for the multi-class classification task, including random forest (RF) [54], SVM (with linear kernel and RBF kernel) [30], Extra Trees (ET) [17], and XGBoost [56]. We used all the 150 privacy policies for training and testing purposes. We used 5-fold cross-validation and 20% as the testing set. Table VII shows the results of all classifiers, indicating that the ET classifier achieved the best performance, with the F1-score reaching 0.882. As can be seen, the performance is generally good enough to support the semi-automated corpus construction process.
TABLE VI: Mappings between sub-label’s format and sub-separator’s format and values in the leading ordinal label descriptor
@@ 5 Others Others
TABLE VII: Results of title and paragraph classifiers
@@
Appendix C Complete result
@@
Table VIII shows the results of 12 type classifiers in both segment and document level evaluation.
TABLE VIII: F1-scores under segment-level and document-level evaluation frameworks (all model types), reported as Macro and Micro averages.
Type Segment-level Document-level
Level-1 nodes All nodes Level-1 nodes All nodes
Macro Micro Macro Micro Macro Micro Macro Micro
Type 1 0.517 0.468 0.453 0.460 0.491 0.452 0.442 0.446
Type 2 0.627 0.606 0.491 0.545 0.561 0.545 0.453 0.501
Type 3 0.564 0.523 0.502 0.506 0.542 0.499 0.480 0.489
Type 4 0.663 0.632 0.535 0.577 0.600 0.574 0.486 0.534
Type 5 0.697 0.689 0.583 0.616 0.657 0.659 0.529 0.598
Type 6 0.736 0.722 0.604 0.643 0.685 0.676 0.553 0.603
Type 7 0.586 0.588 0.412 0.509 0.558 0.564 0.418 0.503
Type 8 0.670 0.674 0.485 0.590 0.637 0.656 0.456 0.575
Type 9 0.621 0.613 0.494 0.552 0.601 0.584 0.472 0.541
Type 10 0.693 0.694 0.531 0.622 0.665 0.658 0.525 0.601
Type 11 0.701 0.685 0.577 0.637 0.659 0.666 0.538 0.606
Type 12 0.740 0.753 0.587 0.684 0.669 0.686 0.542 0.612